OpenStudy (anonymous):

3x^3 over 4z^2 times 12z^4 over 15x^5

OpenStudy (anonymous):

Carry out the multiplication like any other fraction

OpenStudy (anonymous):

how

OpenStudy (anonymous):

\[\frac{ 3x^3 }{ 4z^2 } \times \frac{ 12z^4 }{ 15x^5 }\]

OpenStudy (anonymous):

Multiply 3 and 12 in your numerator and since the x and z are different they will stay the same just multiplied and do the same for the bottom.

OpenStudy (anonymous):

36/60

OpenStudy (anonymous):

So you will arrive at:\[\frac{ (3 \times 12) x^3z^4 }{ (4 \times 15)x^5z^2 }\]

OpenStudy (anonymous):

Correct and from there simplify all your terms.

OpenStudy (anonymous):

how do you simplify the terms

OpenStudy (anonymous):